\begin{thebibliography}{23}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Aguilar et~al.(2020)Aguilar, Kar, and
  Solorio}]{aguilar-etal-2020-lince}
Gustavo Aguilar, Sudipta Kar, and Thamar Solorio. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.lrec-1.223} {{L}in{CE}:
  {A} {C}entralized {B}enchmark for {L}inguistic {C}ode-switching
  {E}valuation}.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 1803--1813, Marseille, France. European Language Resources
  Association.

\bibitem[{Chakravarthi et~al.(2020{\natexlab{a}})Chakravarthi, Jose,
  Suryawanshi, Sherly, and McCrae}]{chakravarthi-etal-2020-sentiment}
Bharathi~Raja Chakravarthi, Navya Jose, Shardul Suryawanshi, Elizabeth Sherly,
  and John~Philip McCrae. 2020{\natexlab{a}}.
\newblock \href {https://www.aclweb.org/anthology/2020.sltu-1.25} {A sentiment
  analysis dataset for code-mixed {M}alayalam-{E}nglish}.
\newblock In \emph{Proceedings of the 1st Joint Workshop on Spoken Language
  Technologies for Under-resourced languages (SLTU) and Collaboration and
  Computing for Under-Resourced Languages (CCURL)}, pages 177--184, Marseille,
  France. European Language Resources association.

\bibitem[{Chakravarthi et~al.(2020{\natexlab{b}})Chakravarthi, Muralidaran,
  Priyadharshini, and McCrae}]{chakravarthi-etal-2020-corpus}
Bharathi~Raja Chakravarthi, Vigneshwaran Muralidaran, Ruba Priyadharshini, and
  John~Philip McCrae. 2020{\natexlab{b}}.
\newblock \href {https://www.aclweb.org/anthology/2020.sltu-1.28} {Corpus
  creation for sentiment analysis in code-mixed {T}amil-{E}nglish text}.
\newblock In \emph{Proceedings of the 1st Joint Workshop on Spoken Language
  Technologies for Under-resourced languages (SLTU) and Collaboration and
  Computing for Under-Resourced Languages (CCURL)}, pages 202--210, Marseille,
  France. European Language Resources association.

\bibitem[{Chakravarthi et~al.(2020{\natexlab{c}})Chakravarthi, Priyadharshini,
  Muralidaran, Suryawanshi, Jose, Sherly, and
  McCrae}]{chakravarthi2020overview}
BR~Chakravarthi, R~Priyadharshini, V~Muralidaran, S~Suryawanshi, N~Jose,
  E~Sherly, and JP~McCrae. 2020{\natexlab{c}}.
\newblock Overview of the track on sentiment analysis for dravidian languages
  in code-mixed text.
\newblock In \emph{Working Notes of the Forum for Information Retrieval
  Evaluation (FIRE 2020). CEUR Workshop Proceedings. In: CEUR-WS. org,
  Hyderabad, India}.

\bibitem[{Chaudhary et~al.(2020)Chaudhary, Raman, Srinivasan, and
  Chen}]{chaudhary2020dictmlm}
Aditi Chaudhary, Karthik Raman, Krishna Srinivasan, and Jiecao Chen. 2020.
\newblock \href {http://arxiv.org/abs/2010.12566} {Dict-mlm: Improved
  multilingual pre-training using bilingual dictionaries}.

\bibitem[{Conneau et~al.(2019)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'a}n, Grave, Ott, Zettlemoyer, and Stoyanov}]{conneau2019unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzm{\'a}n, Edouard Grave, Myle Ott, Luke Zettlemoyer, and
  Veselin Stoyanov. 2019.
\newblock Unsupervised cross-lingual representation learning at scale.
\newblock \emph{arXiv preprint arXiv:1911.02116}.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzmán, Grave, Ott, Zettlemoyer, and Stoyanov}]{Conneau_2020}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzmán, Edouard Grave, Myle Ott, Luke Zettlemoyer, and
  Veselin Stoyanov. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.747} {Unsupervised
  cross-lingual representation learning at scale}.
\newblock \emph{Proceedings of the 58th Annual Meeting of the Association for
  Computational Linguistics}.

\bibitem[{Devlin et~al.(2018)Devlin, Chang, Lee, and
  Toutanova}]{devlin2018bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2018.
\newblock Bert: Pre-training of deep bidirectional transformers for language
  understanding.
\newblock \emph{arXiv preprint arXiv:1810.04805}.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin-etal-2019-bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1423} {{BERT}: Pre-training of
  deep bidirectional transformers for language understanding}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Gupta et~al.(2021)Gupta, Rallabandi, and Black}]{gupta2021task}
Akshat Gupta, Sai~Krishna Rallabandi, and Alan Black. 2021.
\newblock Task-specific pre-training and cross lingual transfer for
  code-switched data.
\newblock \emph{arXiv preprint arXiv:2102.12407}.

\bibitem[{Gupta et~al.(2020)Gupta, Ekbal, and
  Bhattacharyya}]{gupta-etal-2020-semi}
Deepak Gupta, Asif Ekbal, and Pushpak Bhattacharyya. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.findings-emnlp.206} {A
  semi-supervised approach to generate the code-mixed text using pre-trained
  encoder and transfer learning}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EMNLP 2020}, pages 2267--2280, Online. Association for Computational
  Linguistics.

\bibitem[{Gururangan et~al.(2020)Gururangan, Marasović, Swayamdipta, Lo,
  Beltagy, Downey, and Smith}]{Gururangan_2020}
Suchin Gururangan, Ana Marasović, Swabha Swayamdipta, Kyle Lo, Iz~Beltagy,
  Doug Downey, and Noah~A. Smith. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.740} {Don’t stop
  pretraining: Adapt language models to domains and tasks}.
\newblock \emph{Proceedings of the 58th Annual Meeting of the Association for
  Computational Linguistics}.

\bibitem[{Hande et~al.(2020)Hande, Priyadharshini, and
  Chakravarthi}]{hande-etal-2020-kancmd}
Adeep Hande, Ruba Priyadharshini, and Bharathi~Raja Chakravarthi. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.peoples-1.6}
  {{K}an{CMD}: {K}annada {C}ode{M}ixed dataset for sentiment analysis and
  offensive language detection}.
\newblock In \emph{Proceedings of the Third Workshop on Computational Modeling
  of People's Opinions, Personality, and Emotion's in Social Media}, pages
  54--63, Barcelona, Spain (Online). Association for Computational Linguistics.

\bibitem[{Hochreiter and Schmidhuber(1997)}]{doi:10.1162/neco.1997.9.8.1735}
Sepp Hochreiter and Jürgen Schmidhuber. 1997.
\newblock \href {https://doi.org/10.1162/neco.1997.9.8.1735} {Long short-term
  memory}.
\newblock \emph{Neural Computation}, 9(8):1735--1780.

\bibitem[{Khanuja et~al.(2020)Khanuja, Dandapat, Srinivasan, Sitaram, and
  Choudhury}]{khanuja-etal-2020-gluecos}
Simran Khanuja, Sandipan Dandapat, Anirudh Srinivasan, Sunayana Sitaram, and
  Monojit Choudhury. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.acl-main.329}
  {{GLUEC}o{S}: An evaluation benchmark for code-switched {NLP}}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 3575--3585, Online. Association for
  Computational Linguistics.

\bibitem[{Kingma and Ba(2014)}]{kingma2014adam}
Diederik~P. Kingma and Jimmy Ba. 2014.
\newblock \href {http://arxiv.org/abs/1412.6980} {Adam: A method for stochastic
  optimization}.

\bibitem[{Paszke et~al.(2019)Paszke, Gross, Massa, Lerer, Bradbury, Chanan,
  Killeen, Lin, Gimelshein, Antiga, Desmaison, Kopf, Yang, DeVito, Raison,
  Tejani, Chilamkurthy, Steiner, Fang, Bai, and Chintala}]{NEURIPS2019_9015}
Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory
  Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban
  Desmaison, Andreas Kopf, Edward Yang, Zachary DeVito, Martin Raison, Alykhan
  Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu~Fang, Junjie Bai, and Soumith
  Chintala. 2019.
\newblock \href
  {http://papers.neurips.cc/paper/9015-pytorch-an-imperative-style-high-performance-deep-learning-library.pdf}
  {Pytorch: An imperative style, high-performance deep learning library}.
\newblock In H.~Wallach, H.~Larochelle, A.~Beygelzimer, F.~d\textquotesingle
  Alch\'{e}-Buc, E.~Fox, and R.~Garnett, editors, \emph{Advances in Neural
  Information Processing Systems 32}, pages 8024--8035. Curran Associates, Inc.

\bibitem[{Patwa et~al.(2020)Patwa, Aguilar, Kar, Pandey, PYKL, Gamb{\"a}ck,
  Chakraborty, Solorio, and Das}]{patwa2020semeval}
Parth Patwa, Gustavo Aguilar, Sudipta Kar, Suraj Pandey, Srinivas PYKL,
  Bj{\"o}rn Gamb{\"a}ck, Tanmoy Chakraborty, Thamar Solorio, and Amitava Das.
  2020.
\newblock Semeval-2020 task 9: Overview of sentiment analysis of code-mixed
  tweets.
\newblock \emph{arXiv e-prints}, pages arXiv--2008.

\bibitem[{Pires et~al.(2019)Pires, Schlinger, and
  Garrette}]{pires-etal-2019-multilingual}
Telmo Pires, Eva Schlinger, and Dan Garrette. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1493} {How multilingual is
  multilingual {BERT}?}
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 4996--5001, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Sitaram et~al.(2019)Sitaram, Chandu, Rallabandi, and
  Black}]{sitaram2019survey}
Sunayana Sitaram, Khyathi~Raghavi Chandu, Sai~Krishna Rallabandi, and Alan~W
  Black. 2019.
\newblock A survey of code-switched speech and language processing.
\newblock \emph{arXiv preprint arXiv:1904.00784}.

\bibitem[{Wolf et~al.(2020)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac,
  Rault, Louf, Funtowicz, Davison, Shleifer, von Platen, Ma, Jernite, Plu, Xu,
  Le~Scao, Gugger, Drame, Lhoest, and Rush}]{wolf-etal-2020-transformers}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue,
  Anthony Moi, Pierric Cistac, Tim Rault, Remi Louf, Morgan Funtowicz, Joe
  Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien
  Plu, Canwen Xu, Teven Le~Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest,
  and Alexander Rush. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-demos.6} {Transformers:
  State-of-the-art natural language processing}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing: System Demonstrations}, pages 38--45, Online.
  Association for Computational Linguistics.

\bibitem[{Zampieri et~al.(2019)Zampieri, Malmasi, Nakov, Rosenthal, Farra, and
  Kumar}]{zampieri2019semeval}
Marcos Zampieri, Shervin Malmasi, Preslav Nakov, Sara Rosenthal, Noura Farra,
  and Ritesh Kumar. 2019.
\newblock Semeval-2019 task 6: Identifying and categorizing offensive language
  in social media (offenseval).
\newblock \emph{arXiv preprint arXiv:1903.08983}.

\bibitem[{Zampieri et~al.(2020)Zampieri, Nakov, Rosenthal, Atanasova,
  Karadzhov, Mubarak, Derczynski, Pitenis, and
  {\c{C}}{\"o}ltekin}]{zampieri2020semeval}
Marcos Zampieri, Preslav Nakov, Sara Rosenthal, Pepa Atanasova, Georgi
  Karadzhov, Hamdy Mubarak, Leon Derczynski, Zeses Pitenis, and
  {\c{C}}a{\u{g}}r{\i} {\c{C}}{\"o}ltekin. 2020.
\newblock Semeval-2020 task 12: Multilingual offensive language identification
  in social media (offenseval 2020).
\newblock \emph{arXiv preprint arXiv:2006.07235}.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{eacl2021}
\usepackage{times}
\usepackage{latexsym}

\usepackage{multirow}
\usepackage{graphicx}
\usepackage{booktabs}
\usepackage{amsmath}

\usepackage{microtype}

\aclfinalcopy 

\title{SJ\_AJ@DravidianLangTech-EACL2021: Task-Adaptive Pre-Training of Multilingual BERT models for Offensive Language Identification}

\author{Sai Muralidhar Jayanthi \\
    Language Technologies Institute \\
    Carnegie Mellon University \\
  \texttt{sjayanth@cs.cmu.edu} \\\And
  Akshat Gupta \\
  Electrical and Computer Engineering \\
  Carnegie Mellon University \\
  \texttt{akshatgu@andrew.cmu.edu} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
In this paper we present our submission for the EACL 2021-Shared Task on Offensive Language Identification in Dravidian languages. Our final system is an ensemble of mBERT and XLM-RoBERTa models which leverage task-adaptive pre-training of multilingual BERT models with a masked language modeling objective. Our system was ranked 1st for Kannada, 2nd for Malayalam and 3rd for Tamil. \footnote{Code and pretrained models are available at  \href{https://github.com/murali1996/eacl2021-OffensEval-Dravidian}{github.com/murali1996/eacl2021-OffensEval-Dravidian}.}
\end{abstract}

\section{Introduction}
\label{sec:introduction}
The task of offensive language identification aims to identify offensive language used in social media texts, which could be in the form of threats, abusive language, insults \cite{zampieri2020semeval}. Social media platforms like Twitter, Facebook, YouTube are now mediums where a large number of people interact and express their opinions. Conflicts and target abuses are common on social media platform. Offensive language detection is a form of social media monitoring which is best done by building computational models and automatic methods due to the large stream of data involved. 

Code-Switching\footnote{Interchangeably used with the term code-mixing} is a phenomenon common in bilingual and multilingual communities where words from two (or more) languages are used in the same sentence \cite{sitaram2019survey} and is very commonly used for social media interactions. Thus it is vital to build models that are able to process code-switched data for the relevant social media tasks. Previous shared tasks in the domain of offensive language identification have focused on identifying offensive language in English \cite{zampieri2019semeval}, Arabic, Greek, Danish and Turkish \cite{zampieri2020semeval}. The datasets used for this shared task in Dravidian languages are characterized by code-mixing.

Pretrained BERT models \cite{devlin2018bert} and their multilingual versions-- mBERT and XLM-RoBERTa \cite{conneau2019unsupervised} have produced state of the art results in multiple tasks. These models have also been used to built best performing sentiment analysis systems in various shared tasks for code-mixed datasets \cite{patwa2020semeval, chakravarthi2020overview}. In this paper we present our submission for the Shared Task on Offensive Language Identification in Dravidian languages. We employ task-adaptive pre-training with a Masked Language Modeling objective and ensemble predictions from mBERT and XLM-RoBERTa models for our final submission.

\section{Dataset}
\label{sec:dataset}
The dataset statistics for the task of Offensive Language Identification in Malayalam \cite{chakravarthi-etal-2020-sentiment}, Tamil \cite{chakravarthi-etal-2020-corpus} and Kannada \cite{hande-etal-2020-kancmd} are shown in Table \ref{tab:stats1}. The class-wise training dataset statistics are shown in Table \ref{tab:stats2}. \citet{hande-etal-2020-kancmd} outlines a brief about the data collection procedure and some baselines. We see that the datasets are highly imbalanced.\footnote{More details of the task can be found at \href{https://competitions.codalab.org/competitions/27654}{https://competitions.codalab.org/competitions/27654}}

\begin{table}[!htbp]
\small
\hspace{0.5cm}
\resizebox{0.4\textwidth}{!}{%
    \begin{tabular}{c|c|c|c}
    \multirow{2}{*}{Language} & \multicolumn{3}{c}{Split} \\ \cline{2-4}
                              & Train   & Dev    & Test   \\ \hline
     Kannada                   & 6217    & 777    & 778   \\
     Tamil                     & 35139   & 4388   & 4392  \\
    Malayalam                 & 16010   & 1999   & 2001   \\
    
    \end{tabular}
}
\caption{Dataset statistics}
\label{tab:stats1}
\end{table}

\begin{table}[!htbp]
\resizebox{0.47\textwidth}{!}{%
    \begin{tabular}{l|c|c|c}
    \multirow{2}{*}{Class} & \multicolumn{3}{c}{Language} \\ \cline{2-4}
                   & Kannada   & Tamil  & Malayalam \\ \hline
    Not-\{language\} &  1522  & 1454   &  1287  \\ \hline
    Not\_offensive  &  3544  & 25425   &  14153  \\ \hline
    Offensive\_Targeted\_Insult\_Individual &  487  &  2343  &  239  \\
    Offensive\_Targeted\_Insult\_Group &  329  &  2557  &  140  \\
    Offensive\_Targeted\_Insult\_Other &  123  &  454  &  0  \\
    Offensive\_Untargetede &  212  &  2906  & 191   \\ \hline
    Total Training Data & 6217   & 35139   &  16010  \\ \hline
    \end{tabular}
}
\caption{Class-wise training data statistics}
\label{tab:stats2}
\end{table}
\section{Modeling}
\label{sec:modeling}

In this section, we describe the different techniques and models that were used in developing offensive language classifiers.
We use two popular transformer architecture based models- Multilingual BERT  \cite{devlin-etal-2019-bert} and XLM-RoBERTa \cite{Conneau_2020} as our backbone models. \cite{gupta2021task} showed that these multilingual BERT models achieve state of the art performance on Dravidian language code-mixed sentiment analysis datasets. Inspired by some of the recent works on Hindi-English codemixed datasets, such as \citet{khanuja-etal-2020-gluecos} and \citet{aguilar-etal-2020-lince}, we develop solutions based on language model pretraining and transliteration.

\subsection{Architectures}

Motivated by the successes of 
BERT model and its underlying Masked Language Modeling (MLM) pretraining strategy~\cite{devlin-etal-2019-bert}, several 
derivatives evolved over the last couple of years. In this task, since the datasets contain texts in English, Tamil, Kannada and Malayalam in their native scripts as well as in Romanized form, we identify two suitable BERT based models for creating offensive identification classifiers. 

Multilingual-BERT\footnote{\href{https://github.com/google-research/bert/blob/master/multilingual.md}{https://github.com/google-research/bert/multilingual.md}}, aka. mBERT, is based on BERT architecture but pretrained with Wikipedias of 104 different languages including English, Tamil, Kannada and Malayalam. In this work, we specifically use the \textit{cased} version of mBERT\footnote{\href{https://huggingface.co/transformers/multilingual.html?highlight=multilingual\#bert}{https://huggingface.co/transformers/multilingual.html}}. mBERT is pretrained on textual data from different languages with a shared architecture and a shared vocabulary, due to which there is a possible representation sharing across native and transliterated text forms. Corroborating it, \citet{pires-etal-2019-multilingual}
presented its superior zero-shot abilities for an NER task
in Hinglish. 

XLM-RoBERTa, on the other hand, is also a transformer based model but is pretrained on Common Crawl data\footnote{\href{http://data.statmt.org/cc-100/}{http://data.statmt.org/cc-100/}}. The common crawl data consists of textual data both in native scripts as well as in romanized script for some of the languages. Among the three languages in this task, only Tamil has Romanized data in the Common Crawl.

Given a stream of text as input, mBERT splits the text into sub-words by using WordPiece tokenization, whereas XLM-RoBERTa uses SentencePiece tokenization strategy. The sub-tokens are then appended with terminal markers such as \texttt{[CLS]} and \texttt{[SEP]}, indicating the start and end of the input stream, as shown below. 
$$ [CLS], w_1, w_2, ..., w_n, [SEP] $$

Popularized by BERT, the representation corresponding to the \texttt{[CLS]} token in the output layer is utilized for multi-class classification via a softmax layer. We utilize the same methodology in this work. While there are alternative choices for obtaining sentence-level representations, such as \textit{sentence-transformers}\footnote{\href{https://github.com/UKPLab/sentence-transformers}{https://github.com/UKPLab/sentence-transformers}}, we leave their evaluation for future work. We obtain results with this setup and consider them as our baselines.

In an attempt to understand the usefulness of character embeddings when augmented with representations from BERT-based models, we develop a fusion architecture. In this architecture, we first obtain sub-word level representations from mBERT or XLM-RoBERTa models. We then convert them back to word-level representations by averaging the representations of sub-tokens of a given word. If a word is not split into sub-tokens, we use its representation as the word's representations. For every batch of training data, we keep track of words and their corresponding sub-tokens for the purpose of averaging. In a parallel, we also pass character-level embedding sequences through a Bidirectional-LSTM for every word. The hope is that the character-level BiLSTM helps to capture different variations in word patterns due to Romanization, especially useful in the context of social media text.

Once we obtain word-level representations from both character-level BiLSTM as well as BERT-based models, we concatenate the representations at word-level and pass them through a word-level Bidirectional-LSTM. The representation of the last word at the output of BiLSTM is then used for classification.

\begin{table*}[!htbp]
\tiny
\centering
\hspace{0.5cm}
\resizebox{0.8\textwidth}{!}{%

\begin{tabular}{cc|ccc}
\multicolumn{1}{c|}{\multirow{2}{*}{Model}} & \multicolumn{1}{c|}{\multirow{2}{*}{Input Text}} & \multicolumn{3}{c}{F1 / Acc (Dev Split)} \\ \cline{3-5}
\multicolumn{1}{c|}{} & \multicolumn{1}{c|}{} & \multicolumn{1}{c|}{Kannada} & \multicolumn{1}{c|}{Tamil} & \multicolumn{1}{c}{Malayalam} \\
\hline

\multicolumn{5}{c}{Baselines} \\ \hline
mBERT & \multicolumn{1}{c}{\multirow{2}{*}{as-is}} & 67.30 / 70.79 & 76.79 / 79.01 & 95.77 / 96.10 \\
XLM-RoBERTa & \multicolumn{1}{c}{} &  68.70 / 71.56 & 76.28 / 78.19 & 94.12 / 94.85  \\
\hline

\multicolumn{5}{c}{Task-adaptive pretraining} \\  \hline
mBERT & \multicolumn{1}{c}{\multirow{2}{*}{as-is}} & 68.40 / 72.07 & 76.95 / 78.87 & 96.58 / 96.75 \\
XLM-RoBERTa & \multicolumn{1}{c}{} & 69.66 / 72.72 & 77.10 / 78.83 & 96.28 / 96.55 \\
\hline

\multicolumn{5}{c}{Transliteration} \\  \hline
mBERT & \multicolumn{1}{c}{\multirow{2}{*}{romanized}} & 68.84 / 72.07 & 75.16 / 78.49 & 95.72 / 95.80 \\
XLM-RoBERTa & \multicolumn{1}{c}{} & 67.73 / 71.17 & 75.73 / 78.08 & 94.18 / 95.00  \\
\hline

\multicolumn{5}{c}{Fusion Architecture (w/ char-BiLSTM)} \\  \hline
XLM-RoBERTa & \multicolumn{1}{c}{\multirow{1}{*}{as-is}} & 69.79 / 72.97 & 75.51 / 77.87 & 95.78 / 96.05 \\
\hline

\multicolumn{5}{c}{Ensemble (Task-adaptive pretraining))} \\  \hline
mBERT & \multicolumn{1}{c}{\multirow{1}{*}{as-is}} & 70.30 / 74.00 & 76.94 / 79.67 & 96.76 / 97.00 \\
\hline

\end{tabular}

}
\caption{Results on all three languages of the task. \textit{as-is} implies the text is inputted without any preprocessing.}
\label{tab:results1}
\end{table*}

\subsection{Training Methodologies}

In this section, we discuss some techniques to tailor mBERT and XLM-ROBERTa models for the task of offensiveness identification in Dravidian languages.

\paragraph{MLM Pretraining:} 

Both the backbone models are pretrained with their respective corpora using Masked Language Modeling (MLM) objective-- given a sentence, the model randomly masks 15\% of the tokens in the input stream, runs the masked stream through several transformer layers, and then has to predict the masked tokens. Doing so helps the model to learn bi-directional contextual representations, unlike LSTM~\cite{doi:10.1162/neco.1997.9.8.1735} based language modeling. 

Recent works such as \citet{Gururangan_2020} showed that a second phase of pretraining (called domain-adaptive pretraining) with in-domain data can lead to performance gains. Relevant in the realm of codemixing NLP, \citet{khanuja-etal-2020-gluecos} and \citet{aguilar-etal-2020-lince} have shown that such a domain-adaptive pretraining can improve performances for classification tasks in Hinglish. They curated a corpora with millions of entries consisting of codemixed Hinglish in Romanized form. Once the multilingual BERT-based models are trained with such data, they utilized those models for downstream tasks. Their downstream tasks generally contained only texts in Romanized form. \citet{Gururangan_2020} have further shown that adapting the models to the task's unlabeled data (called task-adaptive pretraining) improves performance even after domain-adaptive pretraining.

However, we identify few challenges in order to directly adopt ideologies from the related works to our task at hand. Firstly, the datasets provided for each language in this task consists of texts in English, native script of that language and its transliterated form. Thus, in order to convert everything into Romanized form, we need an accurate transliterator tool. Moreover, even if such a oracle transliterator exists, there could be challenges related to word normalization due to peculiarity of a given language's usage in social media platforms. Secondly, obtaining large amounts of Malayalam-English or Kannada-English code-mixed data for the purpose of pretraining is not readily feasible due to scarcity of relevant public datasets. Thus, we resort to only task-adaptive pretraining. 

Some ideas towards curating corpora for domain-adaptive pretraining could be using machine translation datasets which have fair amounts of code-mixing or by adopting semi-supervised code-mixed data creation techniques \cite{gupta-etal-2020-semi}. Alternate approaches could be developing techniques for efficient representation-sharing between words from native script and their Romanized forms, similar to works such as \citet{chaudhary2020dictmlm}. We leave these explorations to future work.

\paragraph{Transliteration:} In order to evaluate the performance of mBERT and XLM-RoBERTa models when inputted with only Romanized script, we first transliterate the task datasets. To this end, we identify \texttt{indic-trans}\footnote{\href{https://github.com/libindic/indic-trans}{https://github.com/libindic/indic-trans}} as the transliterator. While task-adaptive pretraining could also be conducted with transliterated texts, we leave this exploration to future work.

\begin{table*}[!htbp]
\tiny
\centering
\hspace{0.5cm}
\resizebox{0.8\textwidth}{!}{%

\begin{tabular}{l|c|ccc}
\multicolumn{1}{c|}{\multirow{2}{*}{Model}} & \multicolumn{1}{c|}{\multirow{2}{*}{Input Text}} & \multicolumn{3}{c}{F1 / Acc (Dev Split)} \\ \cline{3-5}
\multicolumn{1}{c|}{} & \multicolumn{1}{c|}{} & \multicolumn{1}{c|}{Kannada} & \multicolumn{1}{c|}{Tamil} & \multicolumn{1}{c}{Malayalam} \\
\hline

mBERT (run-1) & \multicolumn{1}{c|}{\multirow{2}{*}{as-is}} & 69.88 / 73.49 & 76.04 / 78.30 & 96.48 / 96.65 \\
XLM-RoBERTa (run-1) & \multicolumn{1}{c|}{} & 69.72 / 72.59 & 76.80 / 79.40 & 95.80 / 96.25\\ 

mBERT (run-2) & \multicolumn{1}{c|}{\multirow{2}{*}{as-is}} &68.67 / 72.72 & 76.58 / 79.17 & 96.50 / 96.70 \\
XLM-RoBERTa (run-2) & \multicolumn{1}{c|}{}  & 68.39 / 71.69 & 76.20 / 79.03 & 96.30 / 96.70\\

mBERT (run-3) & \multicolumn{1}{c|}{\multirow{2}{*}{as-is}} &66.28 / 70.91 & 76.30 / 79.54 & 96.04 / 96.55 \\
XLM-RoBERTa (run-3) & \multicolumn{1}{c|}{} &69.08 / 71.94 & \textbf{77.86} / 78.71 & 96.17 / 96.30 \\

\hline
Ensemble (Mode) & \multicolumn{1}{c|}{\multirow{1}{*}{as-is}} & \textbf{70.30} / \textbf{74.00} & 76.94 / \textbf{79.67} & \textbf{96.76} / \textbf{97.00} \\
\hline

\end{tabular}

}
\caption{Ensemble results on all three languages. \textit{as-is} implies the text is inputted without any preprocessing.}
\label{tab:results2}
\end{table*}
\section{Experiments}
\label{sec:experiments}

\paragraph{Dataset and Metrics:} The number of training and validation examples for each language are presented in Table~\ref{tab:stats1}. Table~\ref{tab:stats2} presents the class-wise distribution of the training data. As observed, the distribution is skewed with the dominant class being \textit{Not\_offensive}. In this work, we experiment with all three languages presented in the shared task. We use F1 and Accuracy as evaluation metrics\footnote{\href{https://scikit-learn.org/stable/modules/generated/sklearn.metrics.classification_report.html}{sklearn.metrics.classification\_report.html}}.

\paragraph{Implementation details:} For the purpose of training and validation, the corresponding data splits are utilized. Since gold labels for test sets were not previously available, different techniques were compared using evaluations on the validation set itself. For the purpose of task-adaptive pretraining using MLM objective, we combine the train and validation splits of the task data for training BERT models, and utilize test split of the task data for testing MLM pretraining. Due to compute resource limitations, we trim longer sentences to 300 characters (including spaces) and always use a batch size of 8. We do not perform any preprocessing on the text so as to allow for a generalized evaluation across the different datasets in the task.

We use Huggingface library~\cite{wolf-etal-2020-transformers} for implementation of our backbone models. Specifically, we use \texttt{bert-base-multilingual-cased} and \texttt{xlm-roberta-base} models. For experiments related to representation fusion, we implement the BiLSTM models in PyTorch~\cite{NEURIPS2019_9015}. During task specific training for offensive classification, we optimize using the 
BertAdam\footnote{\href{https://github.com/cedrickchee/pytorch-pretrained-BERT/blob/master/pytorch_pretrained_bert/optimization.py}{github.com/cedrickchee/pytorch-pretrained-BERT}}
optimizer for models with a BERT component and with Adam \cite{kingma2014adam} optimizer for the remainder. These optimizers are used with default parameter settings. All experiments are conducted for 5 epochs and we pick the best model using the dev set. We use a hidden size of 128 for character-level BiLSTM and a size of 256 for word-level BiLSTM. We add dropouts with 0.40 rate at the outputs of both BiLSTMs.

\paragraph{Ensemble:} For the final submission, we create a \textit{majority voting} ensemble of 6 models-- 3 of mBERT and 3 of XLM-RoBERTa-- based on task-adaptive pretraining technique. Table~\ref{tab:results2} shows the results of various runs and the ensemble scores across all three languages.

\paragraph{Discussion:} We observe that task-adaptive pretraining improves performance (1-2\% absolute) in case of both the BERT-based models, with major gains obtained for Malayalam dataset (upto 2\% absolute improvement). We also observe that the baseline performances are retained when the input text is fully Romanized, thereby opening up scope for future research along the direction of domain-adaptive pretraining. We also observe that the representation fusion helps in some datasets and degrades performance in others. However, we believe that more evaluation needs to be conducted to ascertain the usefulness of such fusion.

\section{Conclusions}
\label{sec:conslusions}
In this paper, we present our submission for EACL-2021 Shared Task on Offensive Language Identification in Dravidian languages. Our system was an ensemble of 6 different mBERT and XLM-RoBERTa models fine-tuned for the task of offensive language identification. The BERT models were pre-trained on given datasets in Malayalam, Tamil and Kannada. Our system ranked 1st for Kannada, 2nd for Malayalam and 3rd for Kannada. We also proposed a fusion-architecture to leverage character-level, subword-level and word-level embedding to improve performance.


\bibliography{anthology,eacl2021}
\bibliographystyle{acl_natbib}


\end{document}